### v1

# Nonlocal Josephson diode effect in minimal Kitaev chains

Abstract
We study the emergence of the nonlocal Josephson effect in a system composed of three laterally coupled minimal Kitaev chains and exploit it to realize the nonlocal Josephson diode effect. We find that an imbalance between crossed Andreev reflections and electron cotunneling in the middle Kitaev chain gives rise to an asymmetric $2\pi$-periodic phase-dependent Andreev spectrum, controlled by the superconducting phases across the left and right junctions. We then show that the asymmetric Andreev spectrum, formed by hybridized Andreev bound states at the left and right junctions, enables a supercurrent across one junction via the phase difference at the other junction, thereby signaling the nonlocal Josephson effect. Notably, these nonlocal Josephson supercurrents exhibit distinct positive and negative critical currents, demonstrating the realization of the nonlocal Josephson diode effect with highly tunable polarity and efficiencies exceeding 50%. The nonlocal Josephson diode effect requires breaking local time-reversal and local charge-conjugation symmetries, with the latter being unique to minimal Kitaev chains. Our results establish minimal Kitaev chains as a highly controllable platform for engineering nonlocal Josephson phenomena.
Nonlocality is an inherent quantum effect revealing instantaneous action at a distance, tied to the wavefunction nature, and has led to the prediction of several phenomena relevant to quantum technologies [1]. While mostly explored in relation to entanglement [2, 3, 4], the concept of nonlocality has been often studied in superconducting devices. Perhaps the simplest example is a two-terminal superconducting junction, where variations of voltages on one terminal give rise to a differential current on the second terminal, thus determining the so-called nonlocal conductance [5, 6, 7, 8, 9]. More recently, nonlocality was studied in multi-terminal Josephson junctions, where a supercurrent flows across one part of the system due to variations in the superconducting phases elsewhere, defining what is known as the nonlocal Josephson effect (JE) [10]. Since Josephson junctions are key for superconducting circuits and computing [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25], the nonlocal JE is directly relevant for quantum applications while offering a way to further delve the foundations of nonlocality.
The research on nonlocal JE has not only attracted theoretical efforts [10, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35] but its realization has already been reported in many experiments [36, 37, 38, 39, 40, 41, 42]. Most of the studies involved conventional spin-singlet $s$-wave superconductors, leaving unconventional superconductors seldom explored in spite of their intrinsic nonlocal order parameters. One of the simplest systems that hosts controllable unconventional superconductivity with a $p$-wave order parameter can be engineered by connecting two quantum dots (QDs) via a superconductor [43, 44]. This setup represents a minimal Kitaev chain, where electron cotunneling (ECT) and crossed Andreev reflections (CARs) mediate superconducting and normal processes between QDs [43, 45]. Minimal Kitaev chains have lately triggered multiple efforts [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58] and their implementation has already been shown in superconductor-semiconductor hybrids [59, 60, 61, 62]. Despite the significant attention, the nonlocal JE and its impact on minimal Kitaev chains have not yet been studied.
Figure 1: Three minimal Kitaev chains laterally coupled to form a double Josephson junction, where the phase of the left (right) junction is controlled by an external magnetic flux $\Phi_{\rm L(R)}$. Here, a minimal Kitaev chain is formed by connecting two quantum dots (QDs, green) by a superconductor (S, blue). The onsite energy of the QDs is controlled by gates (bars, brown), while the coupling between minimal Kitaev chains can be tuned by an insulating barrier (red).
In this work, we consider three laterally coupled minimal Kitaev chains [Fig. 1] and study the emergence of nonlocal Josephson phenomena. In particular, we demonstrate that, when CAR and ECT in the middle minimal Kitaev chain are distinct, an asymmetric phase-dependent Andreev spectrum appears and is controllable by the superconducting phases across the left and right junctions. We find that this asymmetric Andreev spectrum arises from hybridized Andreev bound states at the left and right junctions and requires breaking local time-reversal and local charge-conjugation symmetries. Notably, the asymmetric Andreev spectrum induces a finite supercurrent across one junction by tuning the superconducting phase difference across the other junction, hence unveiling the nonlocal JE. Even more interesting is that the nonlocal supercurrents develop distinct positive and negative critical currents, giving rise to a nonreciprocal Josephson transport that defines the emergence of a nonlocal Josephson diode effect. We show that this diode effect can achieve robust efficiencies surpassing $50\%$ by controlling CAR and ECT in the middle chain and the phase difference in the second junction. Our findings put forward minimal Kitaev chains for realizing highly tunable nonlocal Josephson phenomena.
JJs formed by three minimal Kitaev chains.—We consider a Josephson system where three minimal Kitaev chains described by
$\begin{split}H_{\alpha}(\phi_{\alpha})&=\varepsilon_{\alpha_{1}}c^{\dagger}_{\alpha_{1}}c_{\alpha_{1}}+\varepsilon_{\alpha_{2}}c^{\dagger}_{\alpha_{2}}c_{\alpha_{2}}\\ &+t_{\alpha}c_{\alpha_{1}}^{\dagger}c_{\alpha_{2}}+\Delta_{\alpha}{\rm e}^{i\phi_{\rm\alpha}}c_{\alpha_{1}}^{\dagger}c_{\alpha_{2}}^{\dagger}+{\rm h.c.}\ \end{split}$ (1)
are laterally coupled [Fig. 1] and modelled by
$\begin{split}H(\phi_{\rm L},\phi_{R})=\sum_{\alpha}H_{\alpha}(\phi_{\alpha})+\tau_{\rm L}c^{\dagger}_{\rm L_{2}}c_{\rm M_{1}}+\tau_{\rm R}c^{\dagger}_{\rm M_{2}}c_{\rm R_{1}}+{\rm h.c.}.\end{split}$ (2)
Here $\alpha={\rm L,M,R}$ labels the left (L), middle (M), and right (R) minimal Kitaev chains formed by two sites $j=1,2$, and $c_{\alpha_{j}}$ ($c_{\alpha_{j}}^{\dagger}$) destroys (creates) an electronic state at site $j=1,2$ in the chain $\alpha$. The first and second terms in $H_{\alpha}(\phi_{\alpha})$ describe onsite energies $\varepsilon_{\alpha_{j}}$ controlled by gates [brown bars in Fig. 1]. The third term represents hopping between sites on the same chain, a characteristic of ECT processes. The fourth term is a spinless superconducting potential, with amplitude $\Delta_{\alpha}$ and phase $\phi_{\alpha}$, describing CAR processes. We model the coupling between the three chains by the second and third terms in Eq. (2) with tunneling amplitudes $\tau_{\rm L(R)}$ across the left (right) junctions. Using a gauge symmetry, we also set the phase of the middle chain as $\phi_{\rm M}=0$, leading to the junction depending on two superconducting phases, $H(\phi_{\rm L},\phi_{\rm R})$. The phase differences $\phi_{\rm L}$ and $\phi_{\rm R}$ between chains can be externally controlled by magnetic fluxes [Fig. 1], as in Refs. [40, 63, 42]. Despite the simplicity, two-site Kitaev chains have already been experimentally realized in QDs coupled by a superconductor-semiconductor system [59, 60, 61, 62, 64], where the control of ECT and CAR was demonstrated [65, 66]. As we show below, the system exhibits a unique phase dependence, enabling a nonlocal JE with nonreciprocal origin.
Breaking and preservation of symmetries.—We begin by analyzing symmetries of the Josephson system in Eq. (2). First of all, the quasi-particle energy derived by Eq. (2) forms a pair (E,E)(-E,E) because of particle-hole symmetry (PHS) of the Bogoliubov-de Gennes Hamiltonian. While the Josephson system respects time-reversal symmetry (TRS), $\mathcal{T}H(\phi_{\rm L},\phi_{\rm R})\mathcal{T}^{-1}=H(-\phi_{\rm L},-\phi_{\rm R})$, where the time-reversal anti-unitary operator $\cal T$} is given by ${\cal T}c_{\alpha_{i}}{\cal T}^{-1}=c_{\alpha_{i}}$, it does not have either $\mathcal{T}H(\phi_{\rm L},\phi_{\rm R})\mathcal{T}^{-1}=H(-\phi_{\rm L},\phi_{\rm R})$ or $\mathcal{T}H(\phi_{\rm L},\phi_{\rm R})\mathcal{T}^{-1}=H(\phi_{\rm L},-\phi_{\rm R})$, implying local TRS is broken either across the left or right junctions, respectively. For $\tau_{\rm L(R)}\neq 0$, the left (right) unction also explicitly breaks the local spatial inversion symmetry, even when the entire system may be spatially inversion-symmetric. Thus, each junction satisfies the standard criterion for the Josephson diode effect [67, 68, 69]. However, our system Eq. (2) has an additional symmetry constraint for the Josephson diode effect, which originates from a special symmetry at the “sweet spot” ($\varepsilon_{\alpha}=0$, $t_{\alpha}=\Delta_{\alpha}$), where the chains support quadratically protected Majorana states [43]. In fact, for $\varepsilon_{\alpha}=0$ and $t_{\alpha}=\Delta_{\alpha}$, each minimal chain has a local charge-conjugation symmetry, ${\cal C}_{\alpha_{1}}H_{\alpha}(\phi_{\alpha}){\cal C}_{\alpha_{1}}^{-1}=H_{\alpha}(-\phi_{\alpha})$, where the non-trivial action of the unitary operator ${\cal C}_{\alpha_{1}}$ is defined by ${\cal C}_{\alpha_{1}}c_{\alpha_{1}}{\cal C}_{\alpha_{1}}^{-1}=c^{\dagger}_{\alpha_{1}}$, ${\cal C}_{\alpha_{1}}c_{\alpha_{2}}{\cal C}_{\alpha_{1}}^{-1}=-e^{i\phi_{\alpha}}c_{\alpha_{2}}$, ${\cal C}_{\alpha_{1}}c_{\beta_{i}}{\cal C}_{\alpha_{1}}^{-1}=c_{\beta_{i}}$ $(\beta\neq\alpha)$. Note that ${\cal C}_{\alpha_{1}}$ switches only an electron at site 1 on chain $\alpha$ with its antiparticle. We also have a similar local charge-conjugation symmetry ${\cal C}_{\alpha_{2}}$, switching only an electron at site 2 with its antiparticle. We can partially keep the symmetries even for the junctions in Eq. (2). For instance, by performing ${\cal C}_{{\rm L}_{2}}$, ${\cal C}_{{\rm M}_{1}}$, and an additional gauge rotation, we have ${\cal C}_{\rm LM}H(\phi_{\rm L},\phi_{\rm R}){\cal C}^{-1}_{\rm LM}=H(-\phi_{\rm L},\phi_{\rm R})$, where the unitary operator ${\cal C}_{\rm LM}$ acts as ${\cal C}_{\rm LM}c_{{\rm L}_{1}}{\cal C}^{-1}_{\rm LM}=e^{i\phi_{\rm L}}c_{{\rm L}_{1}}$, ${\cal C}_{\rm LM}c_{{\rm L}_{2}}{\cal C}^{-1}_{\rm LM}=c^{\dagger}_{{\rm L}_{2}}$, ${\cal C}_{\rm LM}c_{{\rm M}_{1}}{\cal C}_{\rm LM}^{-1}=-c^{\dagger}_{{\rm M}_{1}}$, ${\cal C}_{\rm LM}c_{{\rm M}_{2}}{\cal C}_{\rm LM}^{-1}=c_{{\rm M}_{2}}$, and ${\cal C}_{\rm LM}c_{{\rm R}_{i}}{\cal C}_{\rm LM}^{-1}=c_{{\rm R}_{i}}$. We also have the relation ${\cal C}_{\rm MR}H(\phi_{\rm L},\phi_{\rm R}){\cal C}^{-1}_{\rm MR}=H(\phi_{\rm L},-\phi_{\rm R})$ with an unitary operator ${\cal C}_{\rm MR}$ constructed similarly. These relations provide the same constraint on the quasi-particle spectrum as local inversion, $E_{n}(\phi_{\rm L},\phi_{\rm R})=E_{n}(\pm\phi_{\rm L},\pm\phi_{\rm R})$, and thus, we need to break these local charge-conjugation symmetries, in addition to local TRS, to obtain the Josephson diode effect. For this purpose, we consider $\Delta_{\rm M}\neq t_{\rm M}$ below, which breaks both ${\cal C}_{\rm LM}$ and ${\cal C}_{\rm MR}$.
Figure 2: Energy spectrum as a function of $\phi_{\rm L}$ at distinct $\phi_{\rm R}$ for $t_{\rm M}=\Delta_{\rm M}$ (a-d) and $t_{\rm M}<\Delta_{\rm M}$ (e-h). The vertical dotted line marks $\phi_{\rm L}=\pi$, while the red vertical dashed line marks $\phi_{\rm L}$ at which the dispersing lowest ABSs reach zero energy. Parameters: $\varepsilon_{\alpha}=0$, $t_{\rm L,R}=\Delta_{\rm L,R}\equiv\Delta=1$, $\tau_{\rm L}=1$, $\tau_{\rm R}=0.9$.
Andreev bound states and Andreev molecules.—Having identified the role of the superconducting phase differences $\phi_{\rm L,R}$ on the system symmetries, here we explore how they determine the behavior of the Andreev energy spectrum. This is shown in Fig. 2(a-d) as a function of $\phi_{\rm L}$ when ECT and CAR in the middle chain are the same ($t_{\rm M}=\Delta_{\rm M}$) for distinct values of $\phi_{\rm R}$ at $\Delta_{\rm L,R}=t_{\rm L,R}$, while in Fig. 2(e-h) for ECT weaker than CAR. Since $\Delta_{\rm L,R}=t_{\rm L,R}$, the left (right) chains host zero-energy dispersionless Majorana flat bands located in their first (second) QD, which remain robust against $\phi_{\rm L,R}$ [Fig. 2]. Depending on the ratio between CAR and ECT in the middle chain, the Andreev spectrum exhibits a strong dependence on $\phi_{\rm L,R}$. For $t_{\rm M}=\Delta_{\rm M}$ at $\phi_{\rm R}=0$ [Fig. 2(a)], a pair of dispersionless Andreev bound states (ABSs) appear at energies close to $\pm\Delta$ and intersect a pair of dispersing ABSs near $\phi_{\rm L}=0,2\pi$. See green vertical lines. These dispersionless ABSs originate from the right junction between the middle and right chains, and, as $\phi_{\rm R}$ takes finite values, they hybridize with the dispersing ABSs formed in the left junction between the left and middle chains [Figs. 2(a-d)]. The dispersing hybridized ABSs, which can be seen as Andreev molecules studied before in conventional superconductors [10, 70, 71, 26, 72, 41, 73, 63, 33], reach zero energy at $\phi_{\rm L}=\pi$ (gray vertical line) and become flat bands with $\phi_{\rm L}$ at $\phi_{\rm R}=\pi$ [Fig. 2(c)], always maintaining a symmetric profile about $\phi_{\rm L}=\pi$. When $t_{\rm M}\neq\Delta_{\rm M}$, a nonzero $\phi_{\rm R}$ modifies the phase dependence of the spectrum and, notably, it also makes the spectrum asymmetric about $\phi_{\rm L}=\pi$ (vertical dashed red line), provided $\phi_{\rm R}\neq 2\pi n$ with $n\in\mathbb{Z}$; see Fig. 2(e-h) and Fig. 5 in S1 of the End Matter. This asymmetry moves the zero-energy crossing of the dispersing hybridized ABSs to $\phi_{\rm L}\neq\pi$, which occurs below (above) $\phi_{\rm L}=\pi$ when $\phi_{\rm R}<\pi$ ($\phi_{\rm R}>\pi$) for $t_{\rm M}<\Delta_{\rm M}$ [Fig. 2(e-h)]. The asymmetric spectrum turns out to be gapless over $\phi_{\rm L,R}\in(0,2\pi)$ and remains for $t_{\rm M}>\Delta_{\rm M}$ but with the zero-energy crossing at $\phi_{\rm L}<\pi$ ($\phi_{\rm L}>\pi$) when $\phi_{\rm R}>\pi$ ($\phi_{\rm R}<\pi$) in the latter case [Fig. 5]. A consequence of the asymmetric spectrum under $\phi_{\rm L,R}\neq 0$ and $\Delta_{\rm M}\neq t_{\rm M}$ is that the free energy develops minima away from $\phi_{\rm L}=2\pi n$, unlike conventional Josephson junctions [74, 75]; (See S1 in End Matter). Thus, the imbalance between CAR and ECT in the middle Kitaev chain induces an asymmetric Andreev spectrum at nonzero superconducting phases across the left/right junctions, in line with breaking local TR and local charge-conjugation symmetries.
Figure 3: (a,c,e) Supercurrent as a function of $\phi_{\rm L}$ and $\phi_{\rm R}$ for $t_{\rm M}<\Delta_{\rm M}$ (a), $t_{\rm M}=\Delta_{\rm M}$ (c), and $t_{\rm M}>\Delta_{\rm M}$ (e). (b,d,f) Line cuts as a function of $\phi_{\rm L(R)}$ for distinct $\phi_{\rm R(L)}$. In (d,f), the magenta arrow indicates variations of $\phi_{\rm L,R}$ within $(0,\pi)$ in steps of $0.1\pi$. Parameters: $I_{0}=e\Delta/\hbar$, the rest as in Fig. 2.
Local and nonlocal Josephson effect.—A direct consequence of $\phi_{\rm L,R}\neq 0$ and the imbalance between CAR and ECT in the middle chain is the phase-biased Josephson effect, which we explore here. For this purpose, we numerically obtain the phase-biased Josephson current flowing across the left junction as [76, 77, 24] $I_{\rm L}(\phi_{\rm L},\phi_{\rm R})=(e/\hbar)\sum_{n<0}dE_{n}(\phi_{\rm L},\phi_{\rm R})/d\phi_{\rm L}$  [^1] Similarly, we can obtain the supercurrent flowing across the right junction as $I_{\rm R}(\phi_{\rm L},\phi_{\rm R})=(e/\hbar)\sum_{n<0}dE_{n}(\phi_{\rm L},\phi_{\rm R})/d\phi_{\rm R}$. We have verified that $I_{\rm L,R}(\phi_{\rm L},\phi_{\rm R})$ can also be obtained within a Green’s function approach [102, 77], but obtaining them from the spectrum offers a direct way to identify the contributing states [103, 104, 105, 106, 107, 108, 109].. To visualize the supercurrent, in Figs. 3(a,c,e) we show $I_{\rm L}$ as a function of $\phi_{\rm L}$ and $\phi_{\rm R}$ for distinct cases when CAR in the middle chain is equal (a), weaker (c), and stronger (e) than ECT in the same chain. In Figs. 3(b,d,f) we show $I_{\rm L}$ as a function of $\phi_{\rm L}$ for distinct $\phi_{\rm R}$, while Figs. 3(d,f) also show $I_{\rm L}$ as a function of $\phi_{\rm R}$ for distinct $\phi_{\rm L}$. The first observation is that $I_{\rm L}$ in Figs. 3 strongly depends on $\phi_{\rm L,R}$ and the ratio between ECT ($t_{\rm M}$) and CAR ($\Delta_{\rm M}$) in the middle region, a behavior that originates from the ABSs in Figs. 2. When $\Delta_{\rm M}=t_{\rm M}$, the supercurrent $I_{\rm L}$ is $2\pi$-periodic and symmetric under $\phi_{\rm R}\neq 0$ [Figs. 3(a,b)], having $I_{\rm L}(\phi_{\rm L}=n\pi)=0$ for any $\phi_{\rm R}$ and equal maxima/minima that reduce for $\phi_{\rm R}\neq 0,2\pi$; the free energy in Fig. 6(a) has a minimum at $\phi_{\rm L}$ as for 0-junctions [74, 79]. Also, $I_{\rm L}$ exhibits a sawtooth profile at $\phi_{\rm L}=\pi$ for $\phi_{\rm R}\neq\pi$ due to the zero-energy crossing of the dispersing ABSs [Figs. 2(a-d)], while it smoothens out at $\phi_{\rm R}=\pi$ because the dispersing ABSs become flat [Figs. 2(c)].
At $\Delta_{\rm M}\neq t_{\rm M}$, the supercurrent across the left junction $I_{\rm L}$ develops an intriguing anomalous behavior that varies with $\phi_{\rm L,R}\neq 0$. See Figs. 3(c-f). The most notorious feature is that $\phi_{\rm R}\neq n\pi$ gives rise to $I_{\rm L}(\phi_{\rm L}=n\pi)\neq 0$ with $n\in\mathbb{Z}$ and it is not symmetric about $\phi_{\rm L}=\pi$ anymore, which are consequences of the asymmetric Andreev spectrum shown in Figs. 2(e-h). This implies that, even at $\phi_{\rm L}=0$, a supercurrent across the left junction $I_{\rm L}$ emerges and is controlled by the superconducting phase difference across the right junction $\phi_{\rm R}$, signalling the so-called nonlocal JE which was initially predicted in systems with conventional superconductors [10]. The sawtooth profile of $I_{\rm L}$ also occurs away from $\phi_{\rm L}=\pi$ [Figs. 3(d,f)] due to the shifted zero-energy crossing in the ABSs when $t_{\rm M}\neq\Delta_{\rm M}$. See Figs. 2(e-h) and Figs. 5. It is worth noting that, having $I_{\rm L}(\phi_{\rm L}=0,\phi_{\rm R}\neq n\pi)\neq 0$ in Figs. 3(c-f) also means that a $\phi_{\rm L,0}$-Josephson junction is realized at $\Delta_{\rm M}\neq t_{\rm M}$ [74, 79, 80]. At $\phi_{\rm L}=0$, the current as a function of $\phi_{\rm R}$ can develop a $\pi$-shift for $t_{\rm M}>\Delta_{\rm M}$ [Figs. 3(e,f)], leading to a $\pi$-junction [79]. Moreover, there are also signals of a $\phi$-junction behavior [81, 82, 83, 84], where the current $I_{\rm L}$ passes through zero neither at $\phi_{\rm L}=0,2\pi$ nor $\phi_{\rm L}=\pi$ [Figs. 3(d,f)]. Yet another important effect of $\phi_{\rm L,R}\neq 0$ and $t_{\rm M}\neq\Delta_{\rm M}$ is that the maximum positive and negative values of $I_{\rm L}$ over $\phi_{\rm L}\in(0,2\pi)$ are not the same [Figs. 3(c-f)], producing a nonreciprocity between the maximum currents that can flow across the left junction. As a result, a Josephson setup with three laterally coupled minimal Kitaev chains [Eq. (2)] can realize the nonlocal Josephson effect and nonlocal nonreciprocal Josephson transport only when ECT and CAR in the middle chain are distinct.
Figure 4: (a-c) Critical currents $I_{\rm cL}^{\pm}$ as functions of $\phi_{\rm R}$ for $t_{\rm M}<\Delta_{\rm M}$ (a), $t_{\rm M}=\Delta_{\rm M}$ (b), and $t_{\rm M}>\Delta_{\rm M}$ (c). The dashed and solid lines in (a) correspond to $t_{\rm M}=0.2\Delta_{\rm M}$ and $t_{\rm M}=0.5\Delta_{\rm M}$, while to $t_{\rm M}=1.5\Delta_{\rm M}$ and $t_{\rm M}=1.8\Delta_{\rm M}$ in (c). (d,e) Diode’s quality factor $\eta_{\rm L}$ as a function of $\phi_{\rm R}$ for distinct $t_{\rm M}$ in steps of $0.2\Delta_{\rm M}$, marked by colored bars in (e). (f) $\eta_{\rm L}$ as a function of $\phi_{\rm R}$ for distinct $\varepsilon_{\rm M}$ for $t_{\rm M}<\Delta_{\rm M}$ and $t_{\rm M}>\Delta_{\rm M}$. (g) The same as in (f) but for distinct $\tau_{\rm R}$. Parameters: $I_{0}=e\Delta/\hbar$, the rest as in Fig. 2.
Critical currents and nonlocal Josephson diode effect.—To further explore the nonlocal nonreciprocal Josephson transport, we look at the critical currents across the left junction $I_{\rm cL}^{\pm}(\phi_{\rm R})={\rm max}_{\phi_{\rm L}}[\pm I_{\rm L}(\phi_{\rm L},\phi_{\rm R})]$, where $I_{\rm L}(\phi_{\rm L},\phi_{\rm R})$ is the current flowing across the left junction discussed in the previous section. Figs. 4(a-c) show the critical currents $I_{\rm cL}^{\pm}(\phi_{\rm R})$ as a function of $\phi_{\rm R}$ when ECT in the middle chain is weaker, equal, and stronger than CAR in the same chain. At $t_{\rm M}=\Delta_{\rm M}$, when CAR and ECT in the middle chain are distinct, the positive and negative critical currents are equal ( $I_{\rm cL}^{+}(\phi_{\rm R})=I_{\rm cL}^{-}(\phi_{\rm R})$) and develop a minimum at $\phi_{\rm R}=\pi$ [Fig. 4(b)], leading to the absence of the nonlocal Josephson diode effect. Interestingly, for $t_{\rm M}<\Delta_{\rm M}$, the minima of $I_{\rm cL}^{+}(\phi_{\rm R})$ and $I_{\rm cL}^{-}(\phi_{\rm R})$ are shifted to values below and above $\phi_{\rm R}=\pi$, respectively, giving rise to nonreciprocal phase-biased Josephson currents ($I_{\rm cL}^{+}(\phi_{\rm R})\neq I_{\rm cL}^{-}(\phi_{\rm R})$) flowing across the left junction and fully controlled by $\phi_{\rm R}\neq n\pi$. For $t_{\rm M}>\Delta_{\rm M}$, the shifts in the critical currents is the opposite as for $t_{\rm M}<\Delta_{\rm M}$ but $I_{\rm cL}^{+}(\phi_{\rm R})\neq I_{\rm cL}^{-}(\phi_{\rm R})$ remains. Having $I_{\rm cL}^{+}(\phi_{\rm R})\neq I_{\rm cL}^{-}(\phi_{\rm R})$ defines the nonlocal Josephson diode effect, which requires an imbalance between CAR and ECT in the middle chain and also $\phi_{\rm R}\neq n\pi$. These conditions are consistent with the asymmetric Andreev spectrum [Fig. 2] and with the breaking of local TR and charge-conjugation symmetries discussed below Eq. (2), implying that they are crucial for the nonlocal Josephson diode effect in minimal Kitaev chains.
To characterize the nonlocal diode effect, we analyze its efficiency by studying the quality factor $\eta_{\rm L}(\phi_{\rm R})=[I_{\rm cL}^{+}(\phi_{\rm R})-I_{\rm cL}^{-}(\phi_{\rm R})]/[I_{\rm cL}^{+}(\phi_{\rm R})+I_{\rm cL}^{-}(\phi_{\rm R})]$, where $I_{\rm cL}^{\pm}(\phi_{\rm R})$ are the critical currents across the left junction analyzed in the previous paragraph. We show $\eta_{\rm L}(\phi_{\rm R})$ in Fig.  4(d,e) as a function of $\phi_{\rm R}$ and $t_{\rm M}$, while in Fig.  4(f,g) we inspect the stability of the efficiencies when varying the onsite energies and tunneling across the junction since these conditions are very likely unavoidable in a realistic scenario. In S3 and Fig. 7 of the End Matter, we discuss the robustness of $\eta_{\rm L}(\phi_{\rm R})$ against variations of the onsite energies $\varepsilon_{\alpha}$. In Fig.  4(d,e), we observe that a nonzero quality factor appears when $t_{\rm M}\neq\Delta_{\rm M}$ and $\phi_{\rm R}\neq n\pi$, with $n\in\mathbb{Z}$, and it is enhanced just below and above $\phi_{\rm R}=\pi$ achieving values above $0.5$. While $t_{\rm M}/\Delta_{\rm M}<1$ promotes a quality factor having a sine-line profile with $\phi_{\rm R}$ with a fast sign change across $\phi_{\rm R}=\pi$, the quality factors change sign for $t_{\rm M}/\Delta_{\rm M}>1$, see Fig.  4(d,e). This implies that the diode’s polarity and large efficiencies can be fully controlled nonlocally by $\phi_{\rm R}$ and is thus expected to be more tunable than regular two-terminal Josephson diodes [85, 86, 87, 88, 89, 90, 91, 69, 92, 93, 94, 95, 96, 97, 98, 68, 99, 100, 101]. The large efficiencies of the nonlocal Josephson diode effect are robust against variations of the onsite energies of the middle chain, as demonstrated in Figs. 4(f). Strong variations of the onsite energies in the left and right chains, however, have a detrimental effect although $\eta_{\rm L}(\phi_{\rm R})$ remains robust against weak detuning, see Fig. 7. Moreover, having large tunneling amplitudes favours the large efficiency values of the nonlocal Josephson diode [Fig. 4(g)]. We have also verified that slight deviations from the equal CAR and ECT regime in the left /right chains weakly affects the nonlocal Josephson diode effect.
In conclusion, we have demonstrated the nonlocal Josephson effect and the realization of a nonlocal Josephson diode in a system formed by three laterally coupled minimal Kitaev chains. In particular, we found that when the middle chain exhibits an imbalance between crossed Andreev reflections and electron cotunneling, the phase-dependent Andreev spectrum is asymmetric and can be controlled by the phase difference across the right junction. We have then shown that this asymmetric Andreev spectrum, containing hybridized Andreev bound states from both junctions, produces a supercurrent across the left junction that is controllable by the phase difference across the right junction, hence defining the nonlocal Josephson effect. Notably, we have discovered that these nonlocal Josephson effects enable distinct positive and negative critical currents, giving rise to a nonlocal Josephson diode effect whose polarity is highly tunable and can achieve efficiencies above $50\%$. Given the recent advances on nonlocal Josephson effect [36, 37, 38, 39, 40, 41, 42] and minimal Kitaev chains [59, 60, 61, 62], our findings are very likely within experimental reach.
J. C. acknowledges financial support from the Swedish Research Council (Vetenskapsrådet Grant No. 2021-04121). M. S. was supported by JST CREST Grant No. JPMJCR19T2 and JSPS KAKENHI Grant Nos. JP24K00569 and JP25H01250.